# DALex: Lexicase-like Selection via Diverse Aggregation

### 23 Jan 2024 v1

@@ Abstract
Lexicase selection has been shown to provide advantages over other selection algorithms in several areas of evolutionary computation and machine learning. In its standard form, lexicase selection filters a population or other collection on the basis of randomly ordered training cases that are considered one at a time. This iterated filtering process can be time-consuming, particularly in settings with large numbers of training cases. Such settings include many symbolic regression and deep learning applications. In this paper, we propose a new method that is nearly equivalent to lexicase selection in terms of the individuals that it selects, but which does so significantly more quickly. The new method, called DALex (for Diversely Aggregated Lexicase), selects the best individual with respect to a weighted sum of training case errors, where the weights are randomly sampled. This allows us to formulate the core computation required for selection as matrix multiplication instead of recursive loops of comparisons, which in turn allows us to take advantage of optimized and parallel algorithms designed for matrix multiplication for speedup. Furthermore, we show that we can interpolate between lexicase-like behavior and the behavior of “relaxed” variants of lexicase selection, such as epsilon lexicase selection and batch lexicase selection, by adjusting a single hyperparameter, named “particularity pressure,” which represents the importance granted to each individual training case. Results on program synthesis, deep learning, symbolic regression, and learning classifier systems demonstrate that DALex achieves significant speedups over lexicase selection and its relaxed variants while maintaining almost identical problem-solving performance. Under a fixed computational budget, these savings free up resources that can be directed towards increasing population size or the number of generations, enabling the potential for solving more difficult problems.
@@ 1 Introduction
Genetic algorithms [27] is a subfield of Evolutionary Computation that uses concepts from biological evolution—random variation and survival of the fittest—to solve a wide array of difficult problems. Genetic programming (GP) is a subfield of genetic algorithms that evolves programs or functions that take some inputs and produce some outputs. Common subfields of GP include software synthesis, in which a population of programs evolves to satisfy user-defined training cases, and symbolic regression (SR), in which a population of mathematical expressions evolves to fit a regression dataset. Central to these genetic algorithms is the concept of parent selection, in which members of the current population with desirable characteristics are chosen as the starting point from which to create the next generation of individuals. Lexicase selection [25] and epsilon lexicase selection [33] are state-of-the-art selection algorithms developed for the discrete-error domain and the continuous-error domain, respectively. The key idea behind lexicase selection is that it can be helpful to disaggregate the fitness function, selecting parents by considering training cases one at a time in a random order, all the while filtering out individuals that are not elite relative to the other remaining individuals on the current training case. This selection method has been shown to maintain beneficial diversity in evolved populations [1, 19], especially in terms of “specialists”: individuals who may have a high total error but have a very low error on a subset of the training cases [21, 22]. However, due to its iterative nature, lexicase selection can often take a long time and there is not an obvious way to take advantage of parallelism or single-instruction-multiple-data (SIMD) architectures to speed up its runtime.
@@
In this work, we reexamine the lexicase intuition behind deaggregating the fitness function and we develop an efficient selection method based on randomly aggregating the fitness function at each unbatched selection event. Specifically, we quantify the idea that training cases that occur earlier in a lexicase ordering exert greater selection pressure by taking a randomly weighted average of each individual’s error vector. We formulate this selection mechanism in terms of matrix multiplications, which allows us to take advantage of modern advancements in matrix multiplication algorithms and SIMD architectures to achieve significantly faster runtime compared to lexicase selection. In this work, we distinguish between an unbatched selection event, which takes in the error vectors of the population and return the index of the selected parent, and a batched selection event, which takes in the error vectors of the population and returns an equal number of indices representing the selected parents. For a population of size $n$, creating the next generation would take $n$ unbatched selection events or 1 batched selection event. While we present the theory in terms of an unbatched selection event to simplify the discussion, we implement our selection method and present results using a batched selection event.
@@
This paper is organized as follows: In Section 2, we give a brief overview of lexicase selection and its variants. In Section 3, we describe the Diversely Aggregated Lexicase (DALex) selection method and give a brief theoretical treatment of its properties. In Section 4, we describe our experiments and results in three popular domains in which lexicase selection has had demonstrated success. Empirical results show that DALex replicates the problem-solving success of lexicase selection and its most successful variants while offering significantly reduced runtime.
@@ 2 Background and Related Work
Lexicase selection is a parent selection method that assesses individuals based on each training case in turn, instead of constructing a single scalar fitness value for each individual [25]. During each selection event, the training cases are randomly shuffled. For each training case in the order determined by the random shuffle, candidate individuals in the population that are not “elite” with respect to that training case, i.e., have an error greater than the minimum error on that training case among the remaining candidates, are filtered out. If at any point only a single individual remains, that individual is selected. If all of the training cases are exhausted, then a random individual is selected from those remaining. In other words, lexicase selection chooses an individual based on the first training case, using the rest of the cases in order to break ties. This mechanism has been shown to better maintain population diversity than methods based on aggregated fitness measures [19], which has been shown to improve problem-solving performance [17]. Lexicase selection and its variants have been successfully applied in many diverse problem domains beyond software synthesis, such as learning classifier systems [3], symbolic regression [33], SAT solvers [36], deep learning [11], and evolutionary robotics [44].
@@
In continuous-valued domains like SR, it is often unlikely for more than one individual to share the minimum error on a training case. Therefore, the selection pressure exerted by each training case will be too large, and lexicase selection will not proceed beyond one training case. To combat this, the epsilon-lexicase selection method was developed as a relaxation of lexicase selection [33]. Instead of requiring individuals to have the minimum error out of the current candidates on a training case, epsilon-lexicase filters out individuals with errors exceeding an epsilon threshold above the minimum error. This epsilon threshold is adaptively determined by the population dynamics, specifically the median of the absolute deviations from the median error on each training case.
@@
Batch-lexicase [3, 35, 40] is another lexicase relaxation with improved performance on noisy datasets. Instead of considering each training case in turn, batch-lexicase considers groups of training cases, filtering out individuals whose average accuracy on that group is lower than some threshold. Batch-lexicase has been successfully applied to the field of Learning Classifier Systems (LCS) in Aenugu et al. [3], showing improved performance on noisy datasets and improved generalization with limited training data compared to lexicase selection.
@@
There have been many attempts to speed up lexicase selection. Dolson et al. [12] showed that the exact calculation of lexicase selection probabilities is NP-hard. However, Ding et al. [10] are able to calculate an approximate probability distribution from which to sample individuals. They show that their selection method achieves significantly faster runtime while achieving similar selection frequencies compared to lexicase selection. Ding et al. [9] also propose to use a weighted shuffle of training cases so that more difficult cases are considered first. They show that this technique can significantly reduce the number of training cases considered while suffering minimal degradation of problem-solving power. Batch tournament selection (BTS) [35] orders the training cases by difficulty, then groups them into batches. For each batch of cases, BTS then performs tournament selection once, using the average error on that batch as fitness. Compared to epsilon-lexicase selection, they show competitive problem-solving performance on SR datasets with significantly faster runtime.
@@
Furthermore, many selection methods based on aggregated fitness measures have been proposed for multiobjective genetic algorithms. Fitness sharing is a selection method that attempts to maintain diversity by penalizing individuals for being too close to other individuals in the population, with respect to a user-defined distance metric [8]. On the other hand, implicit fitness sharing does not require a distance metric, and instead scales the reward for each training case by the rest of the population’s performance on that training case before computing the aggregated fitness function [34]. Historically assessed hardness (HAH) is another form of fitness sharing in which errors on each training case are scaled by the population’s historical success rate on that problem [29]. In addition to using a distance metric, NSGA-II sorts individuals into pareto-fronts and conducts tournament selection using the nondomination rank and local crowding factor as fitness values [7].
@@
This work lies at the intersection of the lexicase variants mentioned above but takes a different direction compared to the speedup methods developed so far. For each unbatched selection event, we sample a random weighting of training cases. Then, for each individual, we compute a weighted average of the training case errors and select the individual with the lowest error with respect to the weighted average. In contrast to implicit fitness sharing (IFS) or HAH, we use weights that are randomly chosen, span many more orders of magnitude, and are resampled at each selection event instead of at each generation. In contrast to NSGA-II, this method only selects non-dominated individuals because all of the training case weights are positive after the softmax operation. However, like epsilon and batch lexicase, and in contrast to lexicase selection, this method may select individuals outside of the “Pareto boundary” as defined in La Cava et al. [31]. We formulate this selection method as a vectorized matrix multiplication, selecting all individuals in one go. This allows us to take advantage of modern SIMD architecture and algorithmic advances in matrix multiplication to achieve significantly faster runtime compared to lexicase selection.
3 Diversely Aggregated Lexicase
@@ 3.1 Description
Diversely Aggregated Lexicase selection (DALex) operates by selecting the individual with the lowest average error with respect to a randomly sampled set of weights. It is parameterized by the shape and scale of the distribution from which these weights are sampled. In most of our experiments, we fix the shape to be normal and vary the scale by changing the standard deviation of the distribution. We start with a population of $n$ individuals, where each individual is evaluated on $m$ training cases. We sample an importance score for each training case from the distribution $\mathcal{N}(0,{\tt std})$ where the standard deviation std is a tunable hyperparameter which we call the particularity pressure. Generally speaking, a training case’s importance score quantifies how many times more important it is compared to the rest of the training cases. To obtain the training case weights, we apply the softmax function to the importance scores. This allows us to turn differences between the values of importance scores into differences between the magnitudes of the training case weights. For each individual, we compute its average error w.r.t. the weighting as the dot product of the weight vector and the individual’s error vector. Finally, we select the individual with the lowest average error.
In practice, we conduct all selection events at once, batching multiple weight vectors into a weight matrix and using matrix multiplication in place of the dot product. We also perform an initial “pre-selection” step that differs slightly from other lexicase implementations due to the batched nature of our method: we group individuals into equivalence classes based on their error vectors, select equivalence classes using DALex, and then choose a random individual from each selected class [18].
For a population of $n$ individuals evaluated on $m$ training cases, there will be $k\leq n$ distinct equivalence classes, so the algorithm multiplies a $k\times m$ error matrix with a transposed $n\times m$ weight matrix, giving a worst-case asymptotic runtime of $O(m^{2}n)$, which is the same as that of lexicase selection. However, due to advances in the theory of matrix multiplication, the runtime of DALex can be reduced to $O(n^{\omega(\log_{n}m)})$ using methods such as the Coppersmith-Winograd algorithm [6].
3.2 Motivation
@@
First, we argue that DALex prompts us to reconsider a central tenet of lexicase selection. The success of lexicase selection is commonly attributed to its deaggregation of the fitness function. Aggregated selection methods like tournament or fitness-proportional selection, the reasoning goes, are unable to maintain effective problem solving diversity the same way lexicase selection does. To rationalize our results on DALex, we instead propose the slightly different view that the discrepancy is due to the single unweighted aggregation event preceding these selection methods, which causes a loss of information. Because lexicase selection utilizes the entire high-dimensional error vector instead of simply its mean or sum, it is able to more accurately select promising individuals. Since there are many diverse aggregation events in DALex, it is able to fully utilize the information contained in the individuals’ error vectors, and is therefore able to replicate lexicase selection’s success. Even though each individual selection event operates on a single-dimensional total error, diversity arises from different random weights promoting different individuals in different selection events.
@@
Second, we draw a correspondence between the random orderings in lexicase selection and the random weightings in DALex. Given a random ordering of training cases in lexicase selection, the training cases occurring earlier in the ordering are more relevant to what individual is chosen at that selection event, i.e. exert higher selection pressure. The first training case is paramount, and only if multiple individuals are equally good at the first case do we consider the second case, and so on. In fact, in GP problems, the last case in the ordering is often irrelevant because an individual is usually selected after considering a fraction of the total number of cases. In DALex, the importance scores or training case weights assign explicit values to this notion of selection pressure. Given training case weights of $[1.0,~{}0.01,~{}0.1]$ for example, the first case would be ten times as important as the third case, which in turn is ten times as important as the second case. In the limit of infinite particularity pressure, the differences in importance scores tend towards infinity, so the difference in magnitudes of training case weights tends to infinity, and we recover lexicase selection. As the number of training cases increases or the number of orders of magnitude spanned by the errors on each case increases, the particularity pressure needed to replicate lexicase selection increases. Empirically, we are able to replicate lexicase selection using modest particularity pressures such as 20.
@@
One important result of this interpretation is that it explains why we call the standard deviation hyperparameter the particularity pressure. When importance weights are sampled with high standard deviation, the most important training case will be much more important than the second most, and so on, just like in lexicase selection. As the standard deviation decreases, the importance of a single training case decreases. In other words, high particularity pressure corresponds to increasingly lexicase-like selection [42]. We find that we can achieve similar successes to relaxed forms of lexicase selection such as batch-lexicase or epsilon-lexicase simply by choosing an appropriate particularity pressure. In this sense, DALex unifies the diverse selection methods lexicase, epsilon-lexicase, and batch-lexicase into a single selection method. It is important to note that DALex’s only theoretical guarantee is that of asymptotically lexicase-like selection with infinite particularity pressure. As decreasing the particularity pressure is a different way of relaxation compared to those proposed for epsilon-lexicase and batch-lexicase, we do not expect to be able to replicate the selection dynamics of these lexicase variants to an arbitrary degree of accuracy. However, empirical results show that varying the particularity pressure gives enough flexibility to replicate the successes of these relaxed lexicase variants on the problem domains for which they were designed.

### 09 Feb 2024 v2

# Andrew Ni 11 $0$-0003-3480-4536    Li Ding 22 $0$-0002-1315-1196    Lee Spector 3344 10000-0001-5299-4797

@@ Abstract
Lexicase selection has been shown to provide advantages over other selection algorithms in several areas of evolutionary computation and machine learning. In its standard form, lexicase selection filters a population or other collection based on randomly ordered training cases that are considered one at a time. This iterated filtering process can be time-consuming, particularly in settings with large numbers of training cases, including many symbolic regression and deep learning applications. In this paper, we propose a new method that is nearly equivalent to lexicase selection in terms of the individuals that it selects, but which does so in significantly less time. The new method, called DALex (for Diversely Aggregated Lexicase selection), selects the best individual with respect to a randomly weighted sum of training case errors. This allows us to formulate the core computation required for selection as matrix multiplication instead of recursive loops of comparisons, which in turn allows us to take advantage of optimized and parallel algorithms designed for matrix multiplication for speedup. Furthermore, we show that we can interpolate between the behavior of lexicase selection and its “relaxed” variants, such as epsilon and batch lexicase selection, by adjusting a single hyperparameter, named “particularity pressure,” which represents the importance granted to each individual training case. Results on program synthesis, deep learning, symbolic regression, and learning classifier systems demonstrate that DALex achieves significant speedups over lexicase selection and its relaxed variants while maintaining almost identical problem-solving performance. Under a fixed computational budget, these savings free up resources that can be directed towards increasing population size or the number of generations, enabling the potential for solving more difficult problems.
@@ 1 Introduction
Genetic algorithms [26] is a subfield of evolutionary computation that uses concepts from biological evolution—random variation and fitness-based survival—to solve a wide array of difficult problems. Genetic programming (GP) is a subfield of genetic algorithms that evolves programs or functions that take some inputs and produce some outputs. Common subfields of GP include software synthesis, in which a population of programs evolves to satisfy user-defined training cases, and symbolic regression (SR), in which a population of mathematical expressions evolves to fit a regression dataset. Central to these genetic algorithms is the concept of parent selection, in which members of the current population with desirable characteristics are chosen as the starting point from which to create the next generation of individuals. Lexicase selection [24] and epsilon lexicase selection [32] are state-of-the-art selection algorithms developed for the discrete-error domain and the continuous-error domain, respectively. The key idea behind lexicase selection is that it can be helpful to disaggregate the fitness function, selecting parents by considering training cases one at a time in a random order, all the while filtering out individuals which are not elite relative to the other remaining individuals on the current training case. This selection method has been shown to maintain beneficial diversity in evolved populations [36, 17], especially in terms of “specialists": individuals which may have high total error but have very low errors on a subset of the training cases [19, 20]. However, due to its iterative nature, lexicase selection can often take a long time, and there is not an obvious way to take advantage of parallelism or single-instruction-multiple-data (SIMD) architectures to speed up its runtime.
@@
In this work, we reexamine the lexicase intuition behind disaggregating the fitness function and we develop an efficient selection method based on randomly aggregating the fitness function at each individual selection event. Specifically, we quantify the idea that training cases occurring earlier in a lexicase ordering exert greater selection pressure by taking a randomly weighted average of each individual’s error vector. We formulate this selection mechanism as a matrix multiplication, which allows us to take advantage of modern advancements in matrix multiplication algorithms and SIMD architectures to achieve significantly faster runtime compared to lexicase selection.
@@
This paper is organized as follows: In Section 2, we give a brief overview of lexicase selection and its variants. In Section 3, we describe the Diversely Aggregated Lexicase Selection (DALex) algorithm and give a brief theoretical treatment of its properties. In Section 4, we describe our experiments and results in three popular domains in which lexicase selection has seen demonstrated success. Empirical results show that DALex replicates the problem-solving success of lexicase selection and its most successful variants while offering significantly reduced runtime.
@@ 2 Background and Related Work
Lexicase selection is a parent selection method that assesses individuals based on each training case in turn, instead of constructing a single scalar fitness value for each individual [24]. During each individual selection event, the training cases are randomly shuffled. For each training case in the order determined by the random shuffle, candidate individuals in the population that are not “elite" with respect to that training case, i.e., have an error greater than the minimum error on that training case among the remaining candidates, are filtered out. If at any point only a single individual remains, then that individual is selected. If all of the training cases are exhausted, then a random individual is selected from those individuals still remaining. In other words, lexicase selection chooses an individual based on the first training case, using the rest of the cases in order to break ties. This mechanism has been shown to better maintain population diversity than methods based on aggregated fitness measures [17], which has been shown to improve problem-solving performance [15]. Lexicase selection and its variants have been successfully applied in many diverse problem domains beyond software synthesis, such as learning classifier systems [1], symbolic regression [32], SAT solvers [35], deep learning [9], and evolutionary robotics [44].
@@
In continuous-valued domains like SR, it is often unlikely for more than one individual to share the minimum error on a training case. Therefore, the selection pressure exerted by each training case will be too large, and lexicase selection will not proceed beyond one training case. To combat this, the epsilon-lexicase selection method was developed as a relaxation of lexicase selection [32]. Instead of requiring individuals to have the minimum error out of the current candidates on a training case, epsilon-lexicase selection filters out individuals with errors exceeding an epsilon threshold above the minimum error. This epsilon threshold is adaptively determined by the population dynamics, specifically as the median of the absolute deviations from the median error on each training case.
@@
Batch-lexicase selection [1, 34, 40] is another lexicase relaxation with improved performance on noisy datasets. Instead of considering each training case in turn, batch-lexicase selection considers groups of training cases, filtering out individuals whose average accuracy on that group is lower than some threshold. Batch-lexicase selection has been successfully applied to the field of Learning Classifier Systems (LCS) in Aenugu et al. [1], showing improved performance on noisy datasets compared to lexicase selection.
@@
There have been many attempts to speed up lexicase selection. Dolson et al. [10] showed that the exact calculation of lexicase selection probabilities is NP-hard. However, Ding et al. [8] are able to calculate an approximate probability distribution from which to sample individuals. They show that their selection method achieves significantly faster runtime while achieving similar selection frequencies compared to lexicase selection. Ding et al. [7] also propose to use a weighted shuffle of training cases so that more difficult cases are considered first. They show that this technique can significantly reduce the number of training cases considered while suffering minimal degradation of problem-solving power. Batch tournament selection (BTS) [34] orders the training cases by difficulty, then groups them into batches. BTS then performs tournament selection once per batch of cases, using the average error on that batch as the fitness function. They show that BTS achieves similar problem-solving performance to epsilon lexicase selection on SR datasets with significantly faster runtime.
@@
Furthermore, many selection methods based on aggregated fitness measures have been proposed for multiobjective genetic algorithms. Fitness sharing is a selection method that attempts to maintain diversity by penalizing individuals for being too close to other individuals in the population with respect to a user-defined distance metric [6]. On the other hand, implicit fitness sharing (IFS) does not require a distance metric, and instead scales the reward for each training case by the rest of the population’s performance on that training case before computing the aggregated fitness function [33]. Historically assessed hardness (HAH) is another form of fitness sharing in which errors on each training case are scaled by the population’s historical success rate on that problem [28]. In addition to using a distance metric, NSGA-II sorts individuals into pareto fronts and conducts tournament selection using nondomination rank and local crowding factor as fitness values [5].
@@
This work lies at the intersection of the lexicase variants mentioned above but takes a different direction compared to the speedup methods developed so far. For each individual selection event, we sample a random weighting of training cases. Then, for each individual, we compute a weighted average of the training case errors and select the individual with the lowest error with respect to the weighted average. In contrast to IFS or HAH, we use weights that are randomly chosen, span many more orders of magnitude, and are resampled at each selection event instead of at each generation. In contrast to NSGA-II, this method only selects non-dominated individuals because all of the training case weights are positive after the softmax operation. However, like epsilon and batch lexicase selection, and in contrast to lexicase selection, this method may select individuals outside of the “Pareto boundary" as defined in La Cava et al. [30]. We formulate this selection method as a vectorized matrix multiplication, selecting all individuals in a single, batched selection event. This allows us to take advantage of modern SIMD architecture and algorithmic advances in matrix multiplication to achieve significantly faster runtime compared to lexicase selection.
3 Diversely Aggregated Lexicase Selection
@@ 3.1 Description
Diversely Aggregated Lexicase Selection (DALex) operates by selecting the individual with the lowest average error with respect to a randomly sampled set of weights. It is parameterized by the shape and scale of the distribution from which these weights are sampled. In most of our experiments, we fix the shape to that of the normal distribution and vary the scale by changing the standard deviation of the distribution. For an ablation experiment exploring the effect of different distributions on the performance of DALex, see appendix C. We start with a population of $n$ individuals, where each individual is evaluated on $m$ training cases. We sample an importance score for each training case from the distribution $\mathcal{N}(0,{\tt std})$ where the standard deviation std is a tunable hyperparameter which we call the particularity pressure. Generally speaking, a training case’s importance score quantifies how many times more important it is compared to the rest of the training cases. To obtain the training case weights, we softmax the importance scores. This allows us to turn differences between the values of importance scores into differences between the magnitudes of the training case weights. For each individual, we compute its average weighted error as the dot product of the training case weight vector and the individual’s error vector. Finally, we select the individual with the lowest average weighted error.
In practice, we conduct all individual selection events in one batched selection event, combining multiple weight vectors into a weight matrix and using matrix multiplication in place of the dot product. We report selection runtimes in terms of this batched selection event, i.e. the time taken to go from a population of $n$ individuals to the indices of the $n$ selected parents for the next generation. We also perform an initial “pre-selection” step that differs slightly from other lexicase selection implementations due to the batched nature of our method: we group individuals into equivalence classes based on their error vectors, select equivalence classes using DALex, and then choose a random individual from each selected class [16].
For a population of $n$ individuals evaluated on $m$ training cases, there will be $k\leq n$ distinct equivalence classes, so the algorithm multiplies a $k\times m$ error matrix with a transposed $n\times m$ weight matrix, giving an asymptotic runtime of $O(m^{2}n)$, which is the same as that of lexicase selection. However, due to advances in the theory of matrix multiplication, the runtime of DALex can be reduced to $O(n^{\omega(\log_{n}m)})$ using methods such as the Coppersmith-Winograd algorithm [4].
3.2 Intuition
@@
First, we argue that DALex prompts us to reconsider a central tenet of lexicase selection. The success of lexicase selection is commonly attributed to its disaggregation of the fitness function. As the reasoning goes, aggregated selection methods like tournament or fitness-proportional selection are unable to maintain effective problem solving diversity the same way lexicase selection does. To rationalize our results on DALex, we instead propose the slightly different view that the discrepancy in problem-solving performance is due to the single aggregation event preceding these selection methods, which causes a loss of information. Because lexicase selection utilizes the entire high-dimensional error vector instead of simply the average or sum, it is able to more accurately identify promising individuals. Since DALex uses many diverse aggregation events, it is also able to fully utilize the information contained in the individuals’ error vectors, and is therefore able to replicate the success of lexicase selection. Even though each individual selection event operates on a single-dimensional total error, diversity arises from different random weights promoting different individuals in different selection events.
@@
Second, we draw a correspondence between the random orderings in lexicase selection and the random weightings in DALex. Given a random ordering of training cases in lexicase selection, the cases occurring earlier in the ordering exert more control over which individual is chosen at that selection event, i.e. have higher selection pressure. The first training case is paramount, and only if multiple individuals are equally good at the first case do we consider the second case, and so on. In DALex, the importance scores or training case weights assign explicit values to this notion of selection pressure. Given training case weights of $[1.0,~{}0.01,~{}0.1]$ for example, the first case would be ten times as important as the third case, which in turn is ten times as important as the second case. In the limit of infinite particularity pressure, the differences in importance scores tend towards infinity, so the difference in magnitudes of training case weights tends to infinity, and we recover lexicase selection. As the number of training cases increases or the range of magnitudes spanned by the errors on each case increases, the particularity pressure needed to replicate lexicase selection increases. Empirically, we are able to replicate lexicase selection using modest particularity pressures such as 20.
@@
This interpretation also explains why we call the standard deviation hyperparameter the particularity pressure. When importance weights are sampled with a high standard deviation, the most important training case will be much more important than the second most, and so on, just like in lexicase selection. As the standard deviation decreases, the importance of a single training case decreases. In other words, higher particularity pressures correspond to increasingly lexicase-like selection dynamics[42]. We find that we can achieve similar successes to relaxed forms of lexicase selection such as batch-lexicase or epsilon-lexicase selection simply by choosing an appropriate particularity pressure. In this sense, DALex unifies the diverse selection methods lexicase, epsilon-lexicase, and batch-lexicase selection into a single selection method. It is important to note that DALex’s only theoretical guarantee is that of asymptotically lexicase-like selection with infinite particularity pressure. As decreasing the particularity pressure is a different way of relaxation compared to those proposed for epsilon-lexicase and batch-lexicase selection, we do not expect to be able to replicate the selection dynamics of these lexicase variants to an arbitrary degree of accuracy. However, empirical results show that varying the particularity pressure gives enough flexibility to replicate the successes of these relaxed lexicase variants on the problem domains for which they were designed.